OpenStudy (anonymous):

Help! Write the equation of the line that is perpendicular to the line y=2x+2 and passes through the point (6, 3).

OpenStudy (austinl):

Ok, we need to keep a couple things in mind with this one. The slope of a line that is perpendicular to another is it's negative inverse. What do you think will be the slope of the perpendicular line in this case?

OpenStudy (anonymous):

I really don't understand slope, at all.

OpenStudy (austinl):

Slope is \(\displaystyle \frac{rise}{run}~or~\frac{\Delta y}{\Delta x}\) Make sense? A slope of \(\displaystyle \frac{5}{2}\) would mean that you would go up five units and right 2 units from your y-intercept.

OpenStudy (anonymous):

Got it.

OpenStudy (anonymous):

So, how would I find the slope in this case?

OpenStudy (austinl):

In this case we are given this equation, which is in slope-intercept form, \(y=2x+2\) It's basic form is \(y=mx+b\), where m is slope. What is the slope of this line?

OpenStudy (anonymous):

Oh! 2!

OpenStudy (austinl):

Correct, now we need to find the slope of the line that is perpendicular. Which will be the negative reciprocal. For example, lets say we have a slope of \(\displaystyle \frac{a}{b}\). What would be the negative reciprocal of that?

OpenStudy (anonymous):

negative a/b?

OpenStudy (anonymous):

Or -a/b?

OpenStudy (austinl):

No, a reciprocal of a fraction is when you swap the numerator and the denominator. What do you think now?

OpenStudy (anonymous):

b/a?

OpenStudy (austinl):

\(\displaystyle -\frac{b}{a}\) Now, we have a slope of \(\displaystyle 2~or~\frac{2}{1}\) What do you think that will be?

OpenStudy (anonymous):

What do I think what will be?

OpenStudy (austinl):

The negative reciprocal of the slope from the given problem.

OpenStudy (anonymous):

-1/2?

OpenStudy (austinl):

Correct! So the basic structure of our new line will be, \(\displaystyle y=-\frac{1}{2}+b\) Now to find b, the y-intercept, we need to plug in the points given. \((6,3)\rightarrow(x,y)\) Do you think you could do that? Solve for b with those givens?

OpenStudy (anonymous):

Absolutely, however. Where do I put x? After -1/2?

OpenStudy (austinl):

\((\color{red}{6},\color{blue}{3})\rightarrow(\color{red}{x},\color{blue}{y})\) \(\displaystyle\color{blue}{y}=-\frac{1}{2}\color{red}{x}+b\) Make sense?

OpenStudy (anonymous):

Yes. Thank you so much!

OpenStudy (austinl):

Very cool! Good job!

OpenStudy (austinl):

Ok, to answer your concerns, you PLUG IN THE VALUES into the equation with the +b on the end. And you solve for b and then replace it with it.

OpenStudy (anonymous):

Oh, okay. Thanks!